# Pap2Pat: Towards Automated Paper-to-Patent Drafting using Chunk-based Outline-guided Generation

### v1

@@ Abstract
The patent domain is gaining attention in natural language processing research, offering practical applications in streamlining the patenting process and providing challenging benchmarks for large language models (LLMs). However, the generation of the description sections of patents, which constitute more than 90% of the patent document, has not been studied to date. We address this gap by introducing the task of outline-guided paper-to-patent generation, where an academic paper provides the technical specification of the invention and an outline conveys the desired patent structure. We present Pap2Pat, a new challenging benchmark of 1.8k patent-paper pairs with document outlines, collected using heuristics that reflect typical research lab practices. Our experiments with current open-weight LLMs and outline-guided chunk-based generation show that they can effectively use information from the paper but struggle with repetitions, likely due to the inherent repetitiveness of patent language. We release our data and code.
Data & Codehttps://github.com/boschresearch/Pap2Pat
@@ 1 Introduction
Patenting an invention is time-consuming and requires substantial knowledge of the field of the invention, as well as patent law. Therefore, patent drafting [^1] Formally, one can only draft patent applications, which become patent documents only via an office action, but for short we indistinguishably write patent. is a costly process, motivating measures to increase patent attorneys’ productivity. Some tasks in the patenting process are already technologically supported, e.g., prior art search (Shalaby and Zadrozny 2019; Stamatis 2022; Pujari et al. 2022). In contrast, patent drafting remains a predominantly manual effort. Recently, the interest of studying the capabilities of Large Language Models (LLMs) in the patent domain has increased (Shomee et al. 2024; Jiang and Goetz 2024; Casola and Lavelli 2022; Wang et al. 2024). However, prior work is limited to the generation of the sections Abstract and Claims (Hamborg et al. 2017; Lee 2020; Christofidellis et al. 2022; Lee 2023; Zuo et al. 2024; Lee 2024; Bai et al. 2024). Yet, patents also contain a patent description (comprising sections like Field of the Invention, Background, Summary, and Detailed Description) which constitutes more than 90% of the document. [^2] Measured on our dataset using the Llama-3 tokenizer: (0.7% abstract, 91.8% description, 7.5% claims) To the best of our knowledge, due to the lack of benchmark data, its automatic generation has not been studied to date despite offering both practical and scientific value. Automating this time-intensive drafting process could yield considerable cost savings. From a scientific standpoint, patent descriptions present a significant challenge for LLMs due to their technical complexity, specialized language, and substantial length.
Figure 1: Pap2Pat dataset creation. Patents and papers are matched based on heuristics. Outlines $\text{o}_{i}$ are automatically generated from patent chunks $\text{p}_{i}$ using constrained LLM decoding. Circled numbers correspond to numbers of instances.
The goal of this work is to evaluate LLMs’ proficiency in supporting the drafting of complex patent documents from a given invention specification. [^3] We do not aim to assess their capabilities to autonomously generate new inventions. In practice, the inventor typically drafts an invention report, i.e., a document describing and explaining the invention to the patent attorney as the basis for the patent drafting. In research labs, concurrent academic publishing and patenting is common practice (Murray and Stern 2005), and often the paper itself serves as an invention report. The resulting patent-paper pairs (PPPs) (Murray 2002) have been previously used to study innovation dynamics in the field of economics. While confidential free-form invention reports are not publicly available, scientific publications are. We leverage that fact to build Pap2Pat, a new benchmark of 1.8k carefully matched PPPs. Pap2Pat facilitates the study of LLMs on the complex long-text generation task of patent drafting.
Our approach for creating Pap2Pat, which contains PPPs from various domains, is illustrated in Figure 1. We design heuristics that reliably determine whether a patent and a paper describe the same invention, using metadata like author lists, dates, titles and abstracts. We apply a thorough filtering process to ensure accurate pairings and permissively licensed full texts. For LLM-supported patent drafting, we envision a practical setting in which the patent attorney, given the paper or invention report, creates an outline for the patent. This outline acts as a flexible mechanism to control the document structure and content while keeping manual effort low. This level of control also makes the approach practically usable for patent attorneys, as the effort for post-editing is also kept manageable. With Pap2Pat, we provide automatically generated outlines at varying levels of granularity.
A major challenge of the proposed task is the document length: patent descriptions in Pap2Pat are on average 18k tokens long, some exceeding 180k tokens. We propose chunk-based outline-guided generation, a novel approach to generating long patent documents chunk-by-chunk, guided by a user-defined outline and grounded in the corresponding research paper. We benchmark current open-weight LLMs on this task in both zero-shot and fine-tuning setups. Our main contributions are:
@@
    We create and release Pap2Pat, a new dataset of 1.8k PPPs including meticulously extracted full texts and several outlines for each patent.
@@
    We propose to exploit the parallelism of patents and papers to ground the generation of patent drafts in high-quality technical documents.
@@
    We propose a novel chunk-based outline-guided generation approach to patent drafting.
@@
    We conduct in-depth quantitative and qualitative analyses on Pap2Pat. While our results show encouraging performance, they also highlight limitations of current LLMs and reveal promising directions for future work.

### v3

# Pap2Pat: Benchmarking Outline-Guided Long-Text Patent Generation with Patent-Paper Pairs

@@ Abstract
Dealing with long and highly complex technical text is a challenge for Large Language Models (LLMs), which still have to unfold their potential in supporting expensive and time-intensive processes like patent drafting. Within patents, the description constitutes more than 90% of the document on average. Yet, its automatic generation remains understudied. When drafting patent applications, patent attorneys typically receive invention reports (IRs), which are usually confidential, hindering research on LLM-supported patent drafting. Often, pre-publication research papers serve as IRs. We leverage this duality to build Pap2Pat, an open and realistic benchmark for patent drafting consisting of 1.8k patent-paper pairs describing the same inventions. To address the complex long-document patent generation task, we propose chunk-based outline-guided generation using the research paper as technical specification of the invention. Our extensive evaluation using Pap2Pat and a human case study show that LLMs can effectively leverage information from the paper, but still struggle to provide the necessary level of detail. Fine-tuning leads to more patent-style language, but also to more hallucination. We release our data and code at https://github.com/boschresearch/Pap2Pat.
@@ 1 Introduction
Figure 1: Pap2Pat dataset creation (left) and experimental setup (right).
Securing intellectual property is a long and costly process that requires both deep technical knowledge and expertise in patent law. This motivates the use of technology to boost patent attorney productivity. Natural language processing (NLP) already assists prior art search Shalaby and Zadrozny (2019); Stamatis (2022); Pujari et al. (2021) and patent landscaping (Choi et al., 2022; Pujari et al., 2022). Research on Large Language Models (LLMs) in the patent domain has recently gained momentum Shomee et al. (2024); Jiang and Goetz (2024); Casola and Lavelli (2022); Wang et al. (2024b), but patent drafting remains a largely manual task.
A patent typically consists of claims, which define the invention and the legally relevant scope of protection, and a description, which provides technical details in sections like Field of the Invention, Background, Summary, and Detailed Description. Prior work has primarily focused on generating abstracts and claims Hamborg et al. (2017); Lee (2020); Christofidellis et al. (2022); Lee (2023); Zuo et al. (2024); Lee (2024); Bai et al. (2024b). The description makes up over 90% of the document,  [^1] Measured on our dataset using the Llama-3 tokenizer: 0.7% abstract, 91.8% description, 7.5% claims. implying that large productivity gains are expected from writing support for this section. Yet, generating them remains a significant challenge for LLMs due to their length, technical complexity, and specialized language Wang et al. (2024a, c). Existing work on automatic patent generation suffers from a number of shortcomings, such as ill-posed task setups, lack of open benchmarks, and disregard for patent descriptions (Jiang and Goetz, 2024). In this paper, we tackle all of these issues: we propose a novel setup, evaluation metrics, and outline-guided models for description generation in a realistic setting using open and human-created data.
Inventors typically submit invention reports (IRs), which patent attorneys formalize into patent applications. In many research labs, it is common to use a pre-publication paper as invention report, which leads to so-called patent-paper pairs (PPPs, Murray (2002)), an unrecognized treasure for AI research. To facilitate the study of LLMs on patent drafting, we create Pap2Pat, a new benchmark of 1.8k carefully identified PPPs and patent outlines. We develop and validate a method for reliably matching patents and papers describing the same invention (see Figure 1). For LLM-supported patent drafting, we envision a practical setting in which the attorney, given a paper, provides an outline for the patent. This outline acts as a flexible mechanism to control the document structure and content while keeping manual effort low.
A major challenge of the proposed task is document length: patent descriptions in Pap2Pat are on average 18k tokens long, some exceeding 180k tokens. While current LLMs increasingly support long context windows, they struggle to generate similarly long outputs Liu et al. (2024); Bai et al. (2024a); Wu et al. (2025); Ye et al. (2025). As a remedy, we propose chunk-based outline-guided patent generation (COPGen), which effectively generates long patent documents in chunks.
Evaluating generated patents poses significant challenges due to their length, their technical complexity, and the high cost of manual evaluation. No standard evaluation metrics are established in the literature, and prior work Wang et al. (2024c, a) commonly resorts to standard text similarity metrics which do not work well on long documents Lattimer et al. (2023); Que et al. (2024). We adapt a suite of metrics based on natural language inference (NLI) and authorship attribution to the specific case of evaluating factual correctness, coverage, and language style for long-form patent generation on Pap2Pat. Our main contributions are:
@@
    We create and release Pap2Pat, a new benchmark for patent drafting based on open data that closely aligns with real-world settings.
@@
    We derive a comprehensive suite of automatic evaluation metrics for evaluating generated patents.
@@
    We propose a chunk-based outline-guided patent generation approach with effectively controllable output length. Our method increases coverage considerably while keeping factuality high.
@@
    We conduct extensive evaluations finding that state-of-the-art LLMs can effectively use information from the papers, but still struggle to provide the necessary level of detail, and that fine-tuning leads to much higher stylistic similarity with patents, but substantially decreased factuality.
  5. (5)
    Our human evaluation confirms these findings and indicates promising potential for productivity gains of patent attorneys.